# Empirical Evaluation of Generalizable Automated Program Repair with Large Language Models

### v1

@@ III-D Evaluation Protocol
A challenge in execution-based evaluation of APR is test overfitting [47]: A patch may pass all tests but still fail to implement the intended functionality. To address this issue, the literature distinguishes between plausible patches (those that pass the test suite) and correct patches (those that satisfy the intended requirements). Since a manual review of patches is infeasible at scale, various proxy metrics, such as Syntactic Equivalence with the developer patch (SYE) and Trivial Compiler Equivalence (TCE), where the patch compiles to equivalent bytecode, have been proposed to approximate patch correctness [39]. However, while these metrics display some correlation with correct patch quality, they have not outperformed plausibility consistently as predictors. Ouyang et al. [39] examined the relationship between manually verified correct patches and less expensive metrics (namely, plausibility, TCE, and SYE), and found that TCE and plausibility strongly correlate with patch correctness. A more recent study [48] analyzed test overfitting across several APR approaches and concluded that overfitting may be less problematic than previously assumed. Based on these findings, we adopt plausibility (i.e., the passing of test cases) as our primary evaluation criterion.
@@
Due to the inherent non-determinism of LLM-generated output, results may vary across runs, posing a challenge for achieving reliable, reproducible results [49, 50, 51]. To mitigate the problem, we sample multiple patches per bug from the same LLM, and report the percentage of bugs for which at least one generated patch passes all tests. More precisely, we adopt the approach by Chen et al. [52], generating samples per task and computing an unbiased estimator of as:
where is the number of plausible samples among generations. We report plausible@k for . Based on the standard deviation analysis of plausible@1 for LLM-based APR from [28], we use as a reasonable, yet manageable number of generations. We distributed the generation across three independent runs, each generating five candidate patches using each respective model’s standard setting with a temperature of .
III-E Prompt Design
The goal of this experimental study is to evaluate the automated program repair capabilities of various large language models across different programming languages in a standardized comparison. For this purpose, we use a consistent prompt design across all models and benchmarks. All experiments follow a zero-shot prompt setup, where the prompt does not include any example solutions to avoid potential bias from example quality. Furthermore, the prompts are not iterative [53, 54, 55, 15], meaning that each model processes a sample/query without follow-up interactions. This choice allows us to maintain a comparable scope for the experiments, ensuring that the results reflect the inherent capabilities of the models without being influenced by more complex techniques built on top of them.
@@ III-E Prompt Design
The first prompt is used to evaluate LLMs’ out-of-the-box APR capabilities. It consists of an initial system directive, followed by the code of the buggy method. While previous work has demonstrated that including the failing test case and the resulting error message can improve the bug-fixing performance of LLMs [14, 28], we intentionally adhere to this minimal prompt format to simulate real-world APR scenarios where such test information may be unavailable. The specific prompt is largely adopted from prior work [28], with no prompt engineering or optimization applied to prevent introducing model-specific bias.
@@
The second, more informative prompt incorporates additional information from the failed test cases. In addition to the buggy function, the prompt includes the names and source code of the failing test cases along with the corresponding error messages. Specifically, we adopt the prompt template from RepairBench [42] to ensure comparability.
@@
To investigate the impact of fault localization (FL) on APR performance, we simulate a perfect line-level localization, i.e. the prompt is enriched with explicit hints that highlight the correct location of the buggy code segments requiring modification. In preliminary experiments, we evaluated various methods for integrating such line-level localization information into APR prompts, drawing from prior work [22, 38] as well as developer intuition. The most effective strategy was found to be a simple comment ”TODO: Fix here:” placed immediately before each buggy code snippet. Note that the end of the buggy code segment was not marked. Based on these findings, we adopted this approach for all perfect-FL experiments.
@@
Regarding localization, we also investigate a more realistic setup by drawing the candidate locations from an automated fault localization tool, thus avoiding the assumption of perfect fault localization. In our experiments on Defects4J, we employed the spectrum-based tool FLACOCO [40]. For each defect, the top three suspicious functions identified by FLACOCO were included in the prompt. In instances where a candidate location could not be parsed into a valid function, it was replaced by the next highest-ranked valid function.
IV Results and Analysis
IV-A RQ1: APR Performance of current LLMs across different programming languages
Our first research question aims at comparing LLMs’ language- and prompt-specific APR capabilities. We first present an overall impression across all models and four languages (Section IV-A1) then investigate the influence of adding test information to the prompt (Section IV-A2), and finally address the question whether LLMs tend to perform well on specific languages (Section IV-A3).
@@ IV-A RQ1: APR Performance of current LLMs across different programming languages
| p@ | p@ | p@ | p@ | p@ | p@ | p@ | p@ | p@ | p@ | ||||||
@@
IV-A1 Overall Results with Base Prompt
Table II summarizes the performance of all models across the four benchmarks, reporting plausible@1 (the success rate when generating a single fix) and plausible@5 (success among five generated candidates). Models were evaluated on all reproducible single-hunk bugs using the basic prompt, which contained only the buggy function. The upper section of the table lists closed models, while the lower section presents open models. Most models achieve the best performance on Defects4J, only for DeepSeek R1 (dist.) we see a slightly better performance on BugsPHP with plausible@1 and plausible@5. However, performance varies across languages, and no model consistently outperforms others across all benchmarks. For Java, Claude 3.7 Sonnet performs best with plausible@1, and plausible@5. For Javascript, best results are achieved by Claude 3.5 Haiku with plausible@1 and plausible@5. On BugsPHP, the distilled version of DeepSeek R1 (distilled to Llama 3 70B) achieves the best plausible@1 at and plausible@5 at , while for Python Gemini 2.0 Flash shows the best performance, reaching plausible@1 and plausible@5.
@@ IV-A1 Overall Results with Base Prompt
IV-A2 Influence of Test Information
Next, we compare the base prompt, which contains only the faulty method and a generic repair instruction, with the test prompt, which extends the base prompt to include the names, source code, and error messages of failing test cases.
For this and all subsequent experiments, we broadened the evaluation scope from single-hunk to single-function bugs. To keep the dataset size manageable, we selected a representative subset of 100 bugs per benchmark using stratified sampling over the three complexity levels: We first analyzed the distribution of single-line, single-hunk, and multi-hunk patches per benchmark (see Table I) and sampled accordingly to preserve each benchmark’s original complexity profile. Additionally, we ensured that all repositories included in the original benchmarks were represented in the sampled data. The final distribution is summarized in Table III.
Table IV includes a comparison of the most promising models from Table II for both prompts. Please note that the figures for the base prompt differ from those in Table II, as the evaluation data is different. We observe that adding test information (rows ”Test”) yields substantial improvements compared to the base prompt, randing up to almost % (DeepSeek R1 (dist.) on Python). This confirms results of prior work [14, 28] for the most recent models, and also across multiple programming languages. While all models pick up on the additional test information, the effect appears to be stronger for certain models (particularly for OpenAI’s o3-mini). Notably, the improvements by test information show less variation between the different languages: Across all four benchmarks, test information consistently improves results, with Python (%, averaged over all models) showing the strongest improvements.
@@ IV-A2 Influence of Test Information
| Prompt | Java | JavaScript | PHP | Python | |
| Claude 3.7 Sonnet | Base | ||||
| LL | |||||
| Test | |||||
| Gemini 1.5 Pro | Base | ||||
| LL | |||||
| Test | |||||
| Gemini 2.0 Flash | Base | ||||
| LL | |||||
| Test | |||||
| o3-mini | Base | ||||
| LL | |||||
| Test | |||||
| DeepSeek R1 dist. | Base | ||||
| LL | |||||
| Test | |||||
| Qwen 2.5 Coder | Base | ||||
| LL | |||||
| Test |
IV-A3 Language Specialization
@@
Table II already suggests that model accuracy varies between programming languages, with four different LLMs performing best on the four language-specific benchmarks. Since model complementarity and language specialization are important issues for practitioners, we investigate these questions further. First, the Venn diagrams in Figure 2 displays what model combinations perform well on which subsets of cases. Here, all cases from all benchmarks were joined, and a model is defined to ”solve” an issue if at least one of its generations is plausible. For each model, we find cases that can only be fixed by this particular model. This effect holds for all three prompt types ”basic”, ”test”, and ”line-level localization (LL)” (see Section III-E). We also compute when combining multiple models in a committee, as the chance that at least one of the committee’s models gives a plausible top result:
where refer to the numbers of correct generations by each of the committee’s models. When combining all six models this way, the average chance of repairing a bug increases from % plausible@1 for the single best model (Gemini 1.5 pro) to % for the base prompt.
While such a ”brute-force” committee may be impractical and expensive, smaller, language-specific committees can yield improvements too. For each language, we compare the best model on the respective language with the best two-model committee using the base prompt. We found results to improve for (1) JavaScript from (DeepSeek R1 (dist.)) to % (DeepSeek R1 (dist.) & Claude 3.7), (2) Java from (Gemini 1.5 Pro) to % (Gemini 1.5 Pro & Claude 3.7), (3) PHP from (Gemini 2.0 Flash) to % (Gemini 1.5 Pro & DeepSeek R1 (dist.)), and (4) Python from (Gemini 2.0 Flash) to % (Gemini 2.0 flash & o3-mini). Table V summarizes the results, showing for each benchmark and prompt the best single model and the best two-model committee.
@@
Overall, these results suggest there are substantial differences in LLM repair performance between languages, and that LLMs can complement each other on repair tasks.
| Prompt | Best models | p@1 | p@5 | |
| Java | Base | Gemini 1.5 Pro | ||
| LL | Claude 3.7 | |||
| Test | o3-mini | |||
| Base | Claude 3.7 & Gemini 1.5 Pro | % | % | |
| LL | DeepSeek R1 (dist.) & Claude 3.7 | 37.28% | 48.65% | |
| Test | Claude 3.7 & o3-mini | 59.58% | 69.24% | |
| JavaScript | Base | DeepSeek R1 (dist.) | ||
| LL | Claude 3.7 | |||
| Test | o3-mini | |||
| Base | DeepSeek R1 (dist.) & Claude 3.7 | % | % | |
| LL | DeepSeek R1 (dist.) & Gemini 1.5 Pro | % | % | |
| Test | DeepSeek R1 (dist.) & o3-mini | % | % | |
| Python | Base | Gemini 2.0 | ||
| LL | Gemini 2.0 | |||
| Test | DeepSeek R1 (dist.) | |||
| Base | Gemini 2.0 & o3-mini | % | % | |
| LL | Gemini 2.0 & o3-mini | % | % | |
| Test | Claude 3.7 & DeepSeek R1 (dist.) | % | % | |
| PHP | Base | Gemini 2.0 Flash | ||
| LL | DeepSeek R1 (dist.) | |||
| Test | o3-mini | |||
| Base | DeepSeek R1 (dist.) & Gemini 1.5 Pro | % | % | |
| LL | DeepSeek R1 (dist.) & Gemini 1.5 Pro | % | % | |
| Test | DeepSeek R1 (dist.) & o3-mini | % | % |
IV-B RQ2: Impact of Fault Localization on LLM-based APR
@@
Current work in APR lacks a standardized approach regarding assumptions about fault localization [56]. While end-to-end repair –including fault localization (FL), patch generation, and prioritization – is inherently more challenging than patch generation in isolation, most empirical evaluations of neural program repair assume perfect fault localization as suggested by Liu et al. [57], usually derived directly from developer fixes. This assumption enables objective evaluation but also introduces a disconnect from realistic use cases where fault localization is noisy and imprecise. The impact of this assumption remains largely underexplored, raising concerns about localization bias [58, 57], which may lead to an overestimation of model performance under idealized conditions. Therefore, we investigate two aspects: the impact of FL granularity and the effect of noisy, tool-generated FL on patch generation.
IV-B1 Granularity of Fault Localization

### v2

@@ 3.4. Evaluation Protocol
A challenge in execution-based evaluation of APR is test overfitting (Smith et al., 2015): A patch may pass all tests but still fail to implement the intended functionality. To address this issue, the literature distinguishes between plausible patches (those that pass the test suite) and correct patches (those that satisfy the intended requirements). Since a manual review of patches is infeasible at scale, various proxy metrics, such as Syntactic Equivalence with the developer patch (SYE) and Trivial Compiler Equivalence (TCE), where the patch compiles to equivalent bytecode, have been proposed to approximate patch correctness (Ouyang et al., 2024). However, while these metrics display some correlation with correct patch quality, they have not outperformed plausibility consistently as predictors. Ouyang et al. (Ouyang et al., 2024) examined the relationship between manually verified correct patches and less expensive metrics (namely, plausibility, TCE, and SYE), and found that TCE and plausibility strongly correlate with patch correctness. A more recent study (Petke et al., 2024) analyzed test overfitting across several APR approaches and concluded that overfitting may be less problematic than previously assumed. Based on these findings, we adopt plausibility (i.e., the passing of test cases) as our primary evaluation criterion.
@@
Due to the inherent non-determinism of LLM-generated output, results may vary across runs, posing a challenge for achieving reliable, reproducible results (Ouyang et al., 2023; Sallou et al., 2024; Ouyang et al., 2025). To mitigate the problem, we sample multiple patches per bug from the same LLM, and report the percentage of bugs for which at least one generated patch passes all tests. More precisely, we adopt the approach by Chen et al. (Chen et al., 2021), generating samples per task and computing an unbiased estimator of as:
where is the number of plausible samples among generations. We report pass@k for . Based on the standard deviation analysis of pass@1 for LLM-based APR from (Parasaram et al., 2024), we use as a reasonable, yet manageable number of generations. We distributed the generation across three independent runs, each generating five candidate patches using each respective model’s standard setting with a temperature of .
Significance testing
To assess whether observed differences in between models are statistically meaningful, we apply the Wilcoxon signed-rank test at a significance level of . Results are considered significant when the null hypothesis of equal performance can be rejected. In tables comparing model performance, the best result per language is shown in bold, and all results that are not significantly different from the best are underlined.
3.5. Prompt Design
The goal of this experimental study is to evaluate the automated program repair capabilities of various large language models across different programming languages in a standardized comparison. For this purpose, we use a consistent prompt design across all models and benchmarks. All experiments follow a zero-shot prompt setup, where the prompt does not include any example solutions to avoid potential bias from example quality. Furthermore, the prompts are not iterative (Bouzenia et al., 2024; Zhang et al., 2024a; Xia et al., 2024a; Zhang et al., 2024b), meaning that each model processes a sample/query without follow-up interactions. This choice allows us to maintain a comparable scope for the experiments, ensuring that the results reflect the inherent capabilities of the models without being influenced by more complex techniques built on top of them.
@@ 3.5. Prompt Design
The first prompt is used to evaluate LLMs’ out-of-the-box APR capabilities. It consists of an initial system directive, followed by the code of the buggy method. An example is shown in Listing 1. While previous work has demonstrated that including the failing test case and the resulting error message can improve the bug-fixing performance of LLMs (Xia and Zhang, 2024; Parasaram et al., 2024), we intentionally adhere to this minimal prompt format to simulate real-world APR scenarios where such test information may be unavailable. The specific prompt is largely adopted from prior work (Parasaram et al., 2024), with no prompt engineering or optimization applied to prevent introducing model-specific bias.
@@
The second, more informative prompt incorporates additional information from the failed test cases. It extends the base prompt with the block shown in Listing 2. In addition to the buggy function, the prompt includes the names and source code of the failing test cases along with the corresponding error messages. Specifically, we adopt the prompt template from RepairBench (Silva and Monperrus, 2024) to ensure comparability.
@@
To investigate the impact of fault localization (FL) on APR performance, we simulated perfect line-level localization, i.e. the base prompt is enriched with explicit hints that highlight the correct location of the buggy code segments requiring modification. In preliminary experiments, we evaluated different methods for integrating such line-level localization information into APR prompts, drawing from prior work (Silva et al., 2024; Xia et al., 2023) as well as developer intuition. The most effective strategy was found to be a simple comment ”TODO: Fix here:” placed immediately before each buggy code snippet, as shown in Listing 3. Note that the end of the buggy code segment was not marked. Based on these findings, we adopted this approach for all perfect-FL experiments.
@@
Regarding localization, we also investigate a more realistic setup by drawing the candidate locations from an automated fault localization tool, thus avoiding the assumption of perfect fault localization. In our experiments on Defects4J, we employed the spectrum-based tool FLACOCO (Silva et al., 2021). For each defect, the top three suspicious functions identified by FLACOCO were included in the base prompt, as illustrated in Listing 4. In instances where a candidate location could not be parsed into a valid function, it was replaced by the next highest-ranked valid function. This configuration mirrors practical scenarios where fault localization is imperfect yet still informative.
4. Results and Analysis
4.1. RQ1: APR Performance of current LLMs across different programming languages
Our first research question aims at comparing LLMs’ language- and prompt-specific APR capabilities. We first present an overall impression across all models and four languages (Section 4.1.1), then investigate the influence of adding test information to the prompt (Section 4.1.2), and finally address the question whether LLMs tend to perform well on specific languages (Section 4.1.3).
@@ 4.1. RQ1: APR Performance of current LLMs across different programming languages
| pass@ | pass@ | pass@ | pass@ | pass@ | pass@ | pass@ | pass@ | p@ | pass@ | ||||||
@@
4.1.1. Overall Results with Base Prompt
Table 2 summarizes the performance of all models across the four benchmarks, reporting pass@1 (the success rate when generating a single fix) and pass@5 (success among five generated candidates). Models were evaluated on all reproducible single-hunk bugs using the basic prompt, which contained only the buggy function. The upper section of the table lists closed models, while the lower section presents open models. Most models achieve the best performance on Defects4J, only for DeepSeek R1 (dist.) we see a slightly better performance on BugsPHP with pass@1 and pass@5. However, performance varies across languages, and no model consistently outperforms others across all benchmarks. For Java, Claude 3.7 Sonnet performs best with pass@1, and pass@5. For Javascript, best results are achieved by Claude 3.5 Haiku with pass@1 and pass@5. On BugsPHP, the distilled version of DeepSeek R1 (distilled to Llama 3 70B) achieves the best pass@1 at and pass@5 at , while for Python Gemini 2.0 Flash shows the best performance, reaching pass@1 and pass@5.
@@ 4.1.1. Overall Results with Base Prompt
4.1.2. Influence of Test Information
Next, we compare the base prompt, which contains only the faulty method and a generic repair instruction, with the test prompt, which extends the base prompt to include the names, source code, and error messages of failing test cases. For this and all subsequent experiments, we broadened the evaluation scope from single-hunk to single-function bugs. To keep the dataset size manageable, we selected a representative subset of 100 bugs per benchmark using stratified sampling over the three complexity levels: We first analyzed the distribution of single-line, single-hunk, and multi-hunk patches per benchmark (see Table 1) and sampled accordingly to preserve each benchmark’s original complexity profile. Additionally, we ensured that all repositories included in the original benchmarks were represented in the sampled data. The final distribution is summarized in Table 3.
Table 4 includes a comparison of both prompts for the most promising models from Table 2. These include the four best-performing closed-source models (averaged across all languages) and, among the open models, the best large model and the best medium-sized model. Please note that the figures for the base prompt differ from Table 2, as the evaluation data is different. We observe that adding test information (rows Test) yields substantial improvements compared to the base prompt, reaching up to almost pass@1 (DeepSeek R1 (dist.) on Python). This confirms results of prior work (Xia and Zhang, 2024; Parasaram et al., 2024) for the most recent models, and also across multiple programming languages. While all models pick up on the additional test information, the effect appears to be stronger for certain models (particularly for OpenAI’s o3-mini). Notably, the improvements by test information show less variation between the different languages: Across all four benchmarks, test information consistently improves results, with Python (% pass@1, averaged over all models) showing the strongest improvements.
@@ 4.1.2. Influence of Test Information
| Java | JavaScript | PHP | Python | |||||||||
| Prompt | pass@1 | pass@5 | pass@1 | pass@5 | pass@1 | pass@5 | pass@1 | pass@5 | ||||
| Claude 3.7 Sonnet | Base | |||||||||||
| Test | ||||||||||||
| LL | ||||||||||||
| Gemini 1.5 Pro | Base | |||||||||||
| Test | ||||||||||||
| LL | ||||||||||||
| Gemini 2.0 Flash | Base | |||||||||||
| Test | ||||||||||||
| LL | ||||||||||||
| o3-mini | Base | |||||||||||
| Test | ||||||||||||
| LL | ||||||||||||
| DeepSeek R1 dist. | Base | |||||||||||
| Test | ||||||||||||
| LL | ||||||||||||
| Qwen 2.5 Coder | Base | |||||||||||
| Test | ||||||||||||
| LL | ||||||||||||
4.1.3. Language Specialization
@@
Table 2 has already suggested that model accuracy varies between programming languages, with four different LLMs performing best on the four language-specific benchmarks. Since model complementarity and language specialization are important issues for practitioners, we investigate these issues further. First, the Venn diagrams in Figure 2 displays what model combinations perform well on which subsets of cases. Here, all cases from all benchmarks were joined, and a model is defined to ”solve” an issue if at least one of its generations is plausible. For each model, we find cases that can only be fixed by this particular model. This effect holds for all three prompt types ”basic”, ”test”, and ”line-level localization (LL)” (see Section 3.5). To further investigate how different models complement each other, we evaluate, how changes when the candidates are derived from two models rather than just one. To this end, we calculate when we draw a fraction of of the generations from one model and the remaining generations from a second model:
where represents the number of correct patches among the combined generations. To avoid rounding errors, we approximate this process by simulation: for each , we repeatedly (100 times in our experiments) sample generations from the first model and generations from the second model, and determine as the number of plausible patches in the combined set. We then estimate as the expectation over all simulation runs. Figure 3 shows for ensembles that pair the best-performing model in each benchmark with one of the next four best models, evaluated across different mixing fractions , using the test prompt. In most cases, combining two models improved performance over the best individual model, with more pronounced gains at higher values of .
For instance, with the test prompt improved in 14 out of 16 experiments. Note that, since corresponds to the models’ mixing ratio, the rightmost point in each plot corresponds to sampling only from the best model, and the leftmost point to sampling only from the second model. We found to improve for (1) JavaScript from (o3-mini) to (o3-mini & DeepSeek R1 (dist.)), (2) Java from (o3-mini) to (o3-mini & Claude 3.7), (3) PHP from (o3-mini) to (o3-mini & DeepSeek R1 (dist.)), and (4) Python from (DeepSeek R1 (dist.)) to (Claude 3.7 & DeepSeek R1 (dist.)).
@@
Table 5 summarizes the results, reporting for each benchmark and prompt the best single model and the best two-model ensemble, constructed by combining two-thirds of the generations from the top-performing model with one-third from a complementary model. Overall, the results suggest there are substantial differences in LLM repair performance between programming languages, and demonstrate that LLMs can effectively complement each other on repair tasks.
| Prompt | Best single model | pass@5 | Best ensemble | pass@5 | |||
| Java | Base | Gemini 1.5 Pro | Gemini 1.5 Pro & Claude 3.7 | ||||
| Test | o3-mini | o3-mini & Claude 3.7 | |||||
| LL | Claude 3.7 | Claude 3.7 & DeepSeek R1 (dist.) | |||||
| JavaScript | Base | DeepSeek R1 (dist.) | DeepSeek R1 (dist.) & Claude 3.7 | 21.34% | |||
| Test | o3-mini | o3-mini & DeepSeek R1 (dist.) | |||||
| LL | Claude 3.7 | Claude 3.7 & DeepSeek R1 (dist.) | 35.76% | ||||
| Python | Base | Gemini 2.0 | Gemini 2.0 & Claude 3.7 | 18.39% | |||
| Test | DeepSeek R1 (dist.) | Claude 3.7 & DeepSeek R1 (dist.) | |||||
| LL | Gemini 2.0 | Gemini 2.0 & o3-mini | |||||
| PHP | Base | Gemini 1.5 Pro | Gemini 1.5 Pro & DeepSeek R1 (dist.) | ||||
| Test | o3-mini | o3-mini & DeepSeek R1 (dist.) | |||||
| LL | DeepSeek R1 (dist.) | DeepSeek R1 (dist.) & Gemini 1.5 Pro |
4.2. RQ2: Impact of Fault Localization on LLM-based APR
@@
Current work in APR lacks a standardized approach regarding assumptions about fault localization (Renzullo et al., 2025). While end-to-end repair –including fault localization (FL), patch generation, and prioritization – is inherently more challenging than patch generation in isolation, most empirical evaluations of neural program repair assume perfect fault localization as suggested by Liu et al. (Liu et al., 2019a), usually derived directly from developer fixes. This assumption enables objective evaluation but also introduces a disconnect from realistic use cases where fault localization is noisy and imprecise. The impact of this assumption remains largely underexplored, raising concerns about localization bias (Yang et al., 2017; Liu et al., 2019a), which may lead to an overestimation of model performance under idealized conditions. Therefore, we investigate two aspects: the impact of FL granularity and the effect of noisy, tool-generated FL on patch generation.
4.2.1. Granularity of Fault Localization